# Data augmentation with automated machine learning: approaches and performance comparison with classical data augmentation methods

### v2

@@ Abstract
Data augmentation is arguably the most important regularization technique commonly used to improve generalization performance of machine learning models. It primarily involves the application of appropriate data transformation operations to create new data samples with desired properties. Despite its effectiveness, the process is often challenging because of the time-consuming trial and error procedures for creating and testing different candidate augmentations and their hyperparameters manually. State-of-the-art approaches typically rely on automated machine learning (AutoML) principles. This work presents a comprehensive survey of AutoML-based data augmentation techniques. We discuss various approaches for accomplishing data augmentation with AutoML, including data manipulation, data integration and data synthesis techniques. The focus of this work is on image data augmentation methods. Nonetheless, we cover other data modalities, especially in cases where the specific data augmentations techniques being discussed are more suitable for these other modalities. For instance, since automated data integration methods are more suitable for tabular data, we cover tabular data in the discussion of data integration methods. The work presents extensive discussion of techniques for accomplishing each of the major subtasks of the image data augmentation process: search space design, hyperparameter optimization and model evaluation. Finally, we carried out an extensive comparison and analysis of the performance of automated data augmentation techniques and state-of-the-art methods based on classical augmentation approaches. The results show that AutoML methods for data augmentation currently outperform state-of-the-art techniques based on conventional approaches.
@@ 1.1 Background
Practical implementations of machine learning systems require large data samples to produce satisfactory results. Since data is often not available in sufficient quantities, regularization techniques are critical for achieving good performance. These techniques commonly entail tweaking the machine learning model configuration or applying data augmentation—a range of methods for extending the available data by applying appropriate transformations. The basic idea is to modify training datasets by applying suitable transformations in ways that increase the quantity, representation quality and variability of the datasets. It is used to extend training data in situations where the available samples are insufficient or lack the necessary variability to produce satisfactory performance.
The most commonly used data augmentation techniques include geometric transformations – particularly, rotation, flipping, shearing and scaling–and photometric transformations such as color jittering, solarizaion, brightness, contrast adjustment, noise addition, denoising, and color space conversion. Data augmentation can also involve creating completely new data from scratch [1, 2]. This approach can be useful when the training data for the target application is inaccessible [3]. Methods for synthetic data generation include explicitly creating samples with desired data distribution using computer graphics tools ([2]) or algorithmically generating artificial data with the aid of special deep learning models (e.g., with techniques such as differential neural rendering [4, 5]. Neural style transfer [6], generative modeling techniques such as variational autoencoders (VAEs) [7] and generative adversarial networks (GANs) [8] have also been extensively used to generate synthetic data for training deep learning models. Another way to augment training data is by integrating existing data from several sources (e.g., see [9, 10, 11]). This is a useful way to leverage the large quantities of data available in various forms on the internet and other sources.
@@ 1.2 Limitations of classical data augmentation approaches
Approaches based on GANs are also not guaranteed to produce good results even in cases where sufficiently large and rich datasets are available [15]. Moreover, since generative modeling-based approaches constrain the distribution of generated samples in some way (e.g., by requiring some form of visual similarity of the target domain and generated samples), many potentially useful augmentations that do not satisfy these constrains may be missed. AutoML approaches avoid this limitation by solving the problem in a similarity-independent manner.
@@ 1.4 Motivation for this survey
The concept of automated machine learning (AutoML) has received enormous attention in recent years. One of its most important applications is in data preparation, which includes data pre-processing and data augmentation. Despite there being many surveys on data augmentation methods, very few works have discussed automated data augmentation in sufficient detail. Among the many surveys on AutoML techniques, only a few (e.g., [19, 20, 21, 22, 23]) have discussed the application of AutoML to solve data augmentation problems. Unfortunately, the coverage of automated data augmentation in theses works is very limited in scope and depth. To the best of our knowledge, only Yang et al. [24] and Cheung [25] have presented surveys that are specifically dedicated to automated data augmentation methods. However, these works do not cover many important aspects of AutoML approaches in data augmentation settings, including automated data integration and synthetic data generation methods. Moreover, approaches for the composition of augmentation functions as well as hyperparameter optimization strategies are not given sufficient attention. In addition, to the best of our knowledge, no work has conducted a detailed comparison of the predictive performance of AutoML-based data augmenation approaches and classical methods. This work has been motivated by the emerging importance of AutoML for data augmentation tasks and, as discussed above, the acute lack of coverage of many important issues in the literature.
@@ 1.5 Main contributions
  • We discuss AutoML techniques for solving various data augmentation tasks, including data manipulation, data integration and data synthesis. The last two classes of tasks have not been covered in previous surveys.
@@ 2.2 General procedure for AutoML-based data augmentation
The first step in the AutoML-based data augmentation process is to define of the relevant augmentations to incorporate. A good AutoML data augmentor should incorporate diverse operations that can generate a rich set of data to account for diverse real-world situations. The specification of the types and nature of augmentations typically involves defining different types of primitive image processing operations that transform data in desired ways (e.g., noise perturbation, rotation, scaling, blurring, contrast adjustment, etc.). In addition to the specific transformations, hyperparameters such as the appropriate degree or magnitudes of these transformations (e.g., the range of scale factors, rotation angels, etc.) are also specified. Another aspect of the augmentation problem is to specify how to combine these basic augmentations to compose more complex augmentations, known as augmentation policies. An augmentation policy is generally understood as a set of ordered transformation operations parameterized by selection probabilities and intensity values (see Figure 3). Optimization techniques and heuristic search algorithms are used to find the best augmentation policy. The best augmentation strategy is selected by evaluating the performance of candidate augmentations on the target or a proxy task according to a prior defined performance criterion. This task is generally formulated as a joint optimization problem, where performance on the end task is maximized by jointly optimizing model parameters, augmentation operations and their associated hyperparameters. The most important factors in the choice of search strategy are the accuracy of the resulting policy, the search time and the cost in terms of computational resources (memory and storage).
@@ 3.1 Data manipulation approaches
The easiest and most common way to extend training data is to manipulate existing data by applying appropriate transformations. With this method, typically, a large set of transformation operations, together with possible magnitudes and application probabilities are defined (see Figure 3). There are many ways to generate these basic transformation functions. They can be explicitly specified (e.g., [17, 27, 28] or learned (e.g., [29, 30]). In essence, it consists in designing a set of data transformation operations or implicit parameters or neural models capable of learning transformations on data.
@@ 3.1.1 Image data manipulation
While this approach is widely successful, care must always be exercised in its application; the technique is aimed at changing the very nature of the training data. In particular, the performance of models trained on the resulting data is sensitive to not only the type of transformation operations, but also to the parameters of transformation functions. Consequently, incorrectly specified transformations or aggressive application of operations will often lead to unintended distortions that harms performance. This problem has been reported in several studies, including in the semial works AutoAugment [17] and Population-Based Augmentation [34]. Because AutoML-based data augmentation methods using data manipulation techniques have been extensively covered by other surveys (e.g., [24, 25, 35]), we rather focus on AutoML-based data augmentation strategies that have not received much attention. The most prominent of these approaches are dataset integration and data synthesis methods. These are covered in the next subsections.
@@ 3.1.2 Text data manipulation
Manipulation-based text augmentation creates variations in the original text while still maintaining its meaning and grammatical structure. Approaches for manipulating textual data include rephrasing, randomly deleting or inserting words, replacing worlds by their corresponding synonyms, rearranging sentences in long texts, and introducing noise in the form of typographical errors. Natural language processing techniques [36] are typically used to apply these manipulation methods in AutoML pipelines. Instead of explicitly performing transformations in the AutoML pipeline using basic NLP algorithms, approaches such as Text AutoAugment [37] relies on learning text manipulation polices from input data. More recently, large language models have been seen as a vital tool for automated text augmentation [38, 39]. Besides manipulating text, they can also perform a number of high-level functions such as model configuration or serving as an interface for interaction with the underlying AutoML framework (see [40]).
@@ 3.2.1 Image data integration
Research (e.g., [45, 46, 47]) has shown the effectiveness of mixing different image datasets to overcome the limitations of smaller datasets. It is particularly challenging to integrate image data from different sources as these data often suffer from problems of inconsistent, incomplete, or noisy (i.e., inaccurate) annotations, rendering the effectiveness of automated data integration methods for image modality low. Moreover, data cleaning methods are ill-suited for image modalities. Because of these difficulties, only a few works have attempted using automated machine learning methods to integrate image data from varied sources. Some researchers (e.g., Kim et al. [48, 49]) have proposed to address this challenge by employing dedicated label correction or filtering algorithms in the integration automated data framework.
@@ 3.2.2 Tabular data integration
Techniques for the integration of tabular data [9, 44, 10, 11, 53, 54, 55, 56] generally aim to automatically discover, select and aggregate related data in order to extend a given dataset. Many of the approaches deal with tabular data. Tabular data is typically characterized by problems such as missing values, large discrepancy in data representation, inconsistency of keys, and the presence of a wide variety of variables. This makes the problem of integrating this type of data an extremely challenging but useful task. Chepurko et al. [44] presented an AutoML-based data acquisition strategy aimed at integrating different but related tabular data into a single dataset to provide effective training of machine learning models. Given a set of database tables and corresponding keys, the proposed model automatically searches for, and joins the most related tables in an optimal way. Bai et al. [10] propose neural architecture search (NAS) [57] method to find the best network topology and connections that provide the most effective integration. Kumar et al. [58] demonstrate that integrating relational data from multiple sources does not always lead to performance improvements; it requires a rigorous and careful implementation to achieve improvements. It is the task of the AutoML framework to find integration strategies that provide optimal performance.
@@ 3.2.3 Integrating text and multimodal data
Another challenge with data from different sources may exhibit different statistical distributions [64, 65]. Moreover, in situations where the collected data is from sources created by diverse players with different levels of expertise or conflicting interests and goals, data quality and bias become important issues and need to be handled in the integration process [64]. This is currently achieved using automated cleaning methods [66].
@@ 3.3 Data synthesis with AutoML
Unlike classical data acquisition methods that perform synthesis separately as a distinct subproblem, approaches based on automated machine learning principles typically treat the synthesis task and the target application as a unified problem, where optimal simulation parameters can be jointly learned for the downstream task. This distinction is shown in Figure 5. The workflow of a typical data synthesis method, Task2Sim [73], is shown in Figure 6. Figure 7 shows samples of image data generated by the technique. The approach contrasts sharply with traditional approaches such as A3D [74] that do not make use of feedback signals from the performance on the downstream task to optimize simulation parameters. The synthesis process in these cases, thus, involves extensive manual tuning.
@@ 3.3.2 Tabular data synthesis
Some of the most popular traditional methods for generating synthetic tabular data involve random oversampling CC. This approach is particularly useful for balancing imbalanced data, where minority classes are typically sampled and interpolated to create new datapoints. SMOTE [92], a seminal work among this class of approaches, is commonly used to generate synthetic data to populate tables by means of k-Nearest Neighbor algorithm which interpolates between existing data points to create new samples. These oversampling methods can be incorporated within larger neural frameworks to automate data the synthesis process . For instance, Wang and Pai [93] propose an approach that employs SMOTE together with a GAN model to generate novel tabular clinical data. Similarly, DeepSMOTE [94] combines SMOTE with an encoder-decoder sub-model for synthetic data generation. Aragao et al. [95] propose a self-balancing, synthetic tabular data generation pipeline that utilizes both undersampling and oversampling techniques within an AutoML framework to generate new data. Some new approaches bypass the step of modeling data analytically and instead learn the generation process end-to-end from data. Rashidi et al. [96], for example, devise a so-called Synthetic Tabular Neural Generator (STNG) that trains a dedicated neural network within a larger AutoML framework to specifically generate rich tabular data.
@@ 4.1 Fixed augmentation operations
Many AutoML techniques (e.g., [17, 31, 104, 27, 105, 33, 28, 106] employ a fixed set of pre-defined transformation operations to perform data augmnetation. In these works, the transformation operations themselves are fixed and can only be varied by learning tunable parameters that control hyperparameters such as augmentation intensities and application probabilities. Thus, the task of the AutoML process in this case is reduced to simply learning the most useful combination of these fixed augmentations along with the optimal settings for these hyperparameters. In the procedure, the choice of specific transformation operations and their corresponding magnitudes to include in the search space is based on domain knowledge and intuition.
@@
The process of composing predefined augmentation operations is typically sub-divided into three main sub-problems: (1) composition of transformation operations to be used for augmentation, (2) a specification of (mostly) discrete probabilities with which the transformation operations would be applied on the input data, and (3) a specification of the applicable transformation magnitudes (i.e., augmentation intensities). Again, for simplicity, transformation magnitudes are often discrete values. The original work, AutoAugment (AA) [17], for example, uses a set of 16 common image transformation operations to augment image data for classification tasks. It trains a deep CNN model on each set of transformation operations to learn optimal parameters for effective image augmentations using validation data. Many subsequent works (e.g., [28, 27, 105, 32, 107]) adopt a similar formulation of augmentation operations in the search space as (in) AA [17] but utilize more efficient optimization methods and additional mechanisms to reduce computational requirements. For instance, AdvAA [107] maintained the search space specification in AA [17] but, instead of optimizing the augmentation policy hyperparameters and network parameters separately, jointly optimizes both tasks through adversarial training. Also, instead of RL, a gradient approximation method is used for the policy search.
Figure 10: Augmentation policies of FastAA [105] and the result of applying these policies on sample images of the CIFAR-10 [108] and SVHN [109] datasets.

### v3

@@ Abstract
[…]
Data augmentation is arguably the most important regularization technique commonly used to improve generalization performance of machine learning models. It primarily involves the application of appropriate data transformation operations to create new data samples with desired properties. Despite its effectiveness, the process is often challenging because of the time-consuming trial and error procedures for creating and testing different candidate augmentations and their hyperparameters manually. State-of-the-art approaches are increasingly relying on automated machine learning (AutoML) principles. This work presents a comprehensive survey of AutoML-based data augmentation techniques. We discuss various approaches for accomplishing data augmentation with AutoML, including data manipulation, data integration and data synthesis techniques. The focus of this work is on image data augmentation methods. Nonetheless, we cover other data modalities, especially in cases where the specific data augmentations techniques being discussed are more suitable for these other modalities. For instance, since automated data integration methods are more suitable for tabular data, we cover tabular data in the discussion of data integration methods. The work also presents extensive discussion of techniques for accomplishing each of the major subtasks of the image data augmentation process: search space design, hyperparameter optimization and model evaluation. Finally, we carried out an extensive comparison and analysis of the performance of automated data augmentation techniques and state-of-the-art methods based on classical augmentation approaches. The results show that AutoML methods for data augmentation currently outperform state-of-the-art techniques based on conventional approaches.
@@ 1.1 Background
Practical implementations of machine learning systems require large data samples to produce satisfactory results. Since data is often not available in sufficient quantities, regularization techniques are critical for achieving good performance. These techniques commonly entail tweaking the machine learning model configuration or applying data augmentation—a range of methods for extending the available data by applying appropriate transformations. The basic idea is to modify training datasets by applying suitable transformations in ways that increase the quantity, representation quality and variability of the original data.
The most commonly used data augmentation techniques include geometric transformations – particularly, rotation, flipping, shearing and scaling–and photometric transformations such as color jittering, solarizaion, brightness, contrast adjustment, noise addition, denoising, and color space conversion. Data augmentation can also involve creating completely new data from scratch [1, 2]. This approach can be useful when the training data for the target application is inaccessible [3]. Methods for synthetic data generation include explicitly creating samples with desired data distribution using computer graphics tools ([2]) or algorithmically generating artificial data with the aid of special deep learning models (e.g., with techniques such as differential neural rendering [4, 5]. Neural style transfer [6], generative modeling techniques such as variational autoencoders (VAEs) [7] and generative adversarial networks (GANs) [8] have also been extensively used to generate synthetic data for training deep learning models. Another way to augment training data is by integrating existing data from several sources (e.g., see [9, 10, 11]). This is a useful and a more natural way to augment training data since the approach can leverage the large quantities of data available in various forms on the internet and other sources.
@@ 1.2 Limitations of classical data augmentation approaches
Approaches based on GANs are also not guaranteed to produce good results even in cases where sufficiently large and rich datasets are available [15]. Moreover, since generative modeling-based approaches constrain the distribution of generated samples in some way (e.g., by requiring some form of similarity of the target domain and generated samples), many potentially useful augmentations that do not satisfy these constrains may be missed. AutoML approaches avoid this limitation by solving the problem in a similarity-independent manner.
@@ 1.4 Motivation for this survey
The concept of automated machine learning (AutoML) has received enormous attention in recent years. One of its most important applications is in data preparation, which includes data pre-processing and data augmentation. Despite there being many surveys on data augmentation methods [19, 20, 21], very few works have discussed automated data augmentation in sufficient detail. Among the many surveys on AutoML techniques, only a few (e.g., [22, 23, 24, 25, 26]) have discussed the application of AutoML to solve data augmentation problems. Unfortunately, the coverage of automated data augmentation in theses works is very limited in scope and depth. To the best of our knowledge, only Yang et al. [27] and Cheung [28] have presented surveys that are specifically dedicated to automated data augmentation methods. However, these works do not cover many important aspects of AutoML approaches in data augmentation settings, including automated data integration and synthetic data generation methods. Moreover, approaches for the composition of augmentation functions as well as hyperparameter optimization strategies are not given sufficient attention. In addition, to the best of our knowledge, no work has conducted a detailed comparison of the predictive performance of AutoML-based data augmenation approaches and classical methods. This work has been motivated by the emerging importance of AutoML for data augmentation tasks and, as discussed above, the acute lack of coverage of many important issues in the literature.
@@ 1.5 Main contributions
  • We discuss AutoML techniques for performing various data augmentation tasks, including data manipulation, data integration and data synthesis. The last two classes of tasks have not been covered in previous surveys.
@@ 2.2 General procedure for AutoML-based data augmentation
The first step in the AutoML-based data augmentation process is to define the relevant augmentations to incorporate. A good AutoML data augmentor should incorporate diverse operations that can generate a rich set of data to account for diverse real-world situations. The specification of the types and nature of augmentations typically involves defining different types of primitive data processing operations that transform data in desired ways (e.g., noise perturbation, rotation, scaling, blurring, contrast adjustment, etc.). In addition to the specific transformations, hyperparameters such as the degree or magnitudes of these transformations (e.g., the range of scale factors, rotation angels, etc.) are also specified. Another aspect of the augmentation problem is to specify how to combine these basic augmentations to compose more complex augmentations, known as augmentation policies. An augmentation policy is generally understood as a set of ordered transformation operations parameterized by selection probabilities and intensity values (see Figure 3). Optimization techniques and heuristic search algorithms are used to find the best augmentation policy. The best augmentation strategy is selected by evaluating the performance of candidate augmentations on the target or a proxy task according to a prior defined performance criterion. This task is generally formulated as a joint optimization problem, where performance on the end task is maximized by jointly optimizing model parameters, augmentation operations and their associated hyperparameters. The most important factors in the choice of search strategy are the accuracy of the resulting policy, the search time and the cost in terms of computational resources.
@@ 3.1 Data manipulation approaches
The easiest and most common way to extend training data is to manipulate existing data by applying appropriate transformations. With this method, typically, a large set of transformation operations, together with possible magnitudes and application probabilities are defined (see Figure 3). There are many ways to generate these basic transformation functions. They can be explicitly specified (e.g., [17, 30, 31] or learned (e.g., [32, 33]). In essence, this approach consists in designing a set of data transformation operations or implicit parameters or neural models capable of learning transformations on data.
@@ 3.1.1 Image data manipulation
While this approach is widely successful, care must always be exercised in its application; the technique is aimed at changing the very nature of the training data. In particular, the performance of models trained on the resulting data is sensitive to not only the type of transformation operations, but also to the parameters of transformation functions. Consequently, incorrectly specified transformations or aggressive application of operations will often lead to unintended distortions that harm performance. This problem has been reported in several studies, including in the semial works AutoAugment [17] and Population-Based Augmentation [37]. Because AutoML-based data augmentation methods based on data manipulation techniques have been extensively covered by other surveys (e.g., [27, 28, 38]), we rather focus on AutoML-based data augmentation strategies that have not received much attention. The most prominent of these approaches are dataset integration and data synthesis methods. These are covered in the next subsections.
@@ 3.1.2 Text data manipulation
Manipulation-based text augmentation creates variations in the original text while still maintaining the meaning and grammatical structure. Approaches for manipulating textual data include rephrasing, randomly deleting or inserting words, replacing words by their corresponding synonyms, rearranging sentences in long texts, and introducing noise in the form of typographical errors. Natural language processing techniques [39] are typically used to apply these manipulation methods in AutoML pipelines. Instead of explicitly performing transformations in the AutoML pipeline using basic NLP algorithms, approaches such as Text AutoAugment [40] relies on learning text manipulation polices from input data. More recently, large language models have been seen as a vital tool for automated text augmentation [41, 42]. Besides manipulating text, they can also perform a number of high-level functions such as model configuration or serving as an interface for interaction with the underlying AutoML framework (see [43]).
@@ 3.2.1 Image data integration
Research (e.g., [48, 49, 50]) has shown the effectiveness of mixing different image datasets to overcome the limitations of smaller datasets. Despite the attractiveness of this method, it is particularly challenging to integrate image data from different sources as these data often suffer from problems of inconsistent, incomplete, or noisy (i.e., inaccurate) annotations, rendering the effectiveness of automated data integration methods for image modality low. Moreover, data cleaning methods are ill-suited for image modalities. Because of these difficulties, only a few works have attempted using automated machine learning methods to integrate image data from varied sources. Some researchers (e.g., Kim et al. [51, 52]) have proposed to address this challenge by employing dedicated label correction or filtering algorithms in the integration framework.
@@ 3.2.2 Tabular data integration
Techniques for the integration of tabular data [9, 47, 10, 11, 56, 57, 58, 59] generally aim to automatically discover, select and aggregate related data in order to extend a given dataset. Tabular data is typically characterized by problems such as missing values, large discrepancy in data representation, inconsistency of keys, and the presence of a wide variety of variables. This makes the problem of integrating this type of data an extremely challenging but useful task. Chepurko et al. [47] presented an AutoML-based data acquisition strategy aimed at integrating different but related tabular data into a single dataset to provide effective training of machine learning models. Given a set of database tables and corresponding keys, the proposed model automatically searches for, and joins the most related tables in an optimal way. Bai et al. [10] propose neural architecture search (NAS) [60] method to find the best network topology and connections that provide the most effective integration. Kumar et al. [61] demonstrate that integrating relational data from multiple sources does not always lead to performance improvements; it requires a rigorous and careful implementation to achieve improvements. It is the task of the AutoML framework to find integration strategies that provide optimal performance.
@@ 3.2.3 Integrating text and multimodal data
Another challenge is that data from different sources may exhibit different statistical distributions [67, 68]. Moreover, in situations where the collected data is from sources created by diverse players with different levels of expertise or conflicting interests and goals, data quality and bias become important issues and need to be handled in the integration process [67]. This is currently achieved using automated cleaning methods [69].
@@ 3.3 Data synthesis with AutoML
Unlike classical data acquisition methods that perform synthesis separately as a distinct subproblem, approaches based on automated machine learning principles typically treat the synthesis task and the target application as a unified problem, where optimal simulation parameters can be jointly learned for the downstream task. This distinction is shown in Figure 5. The workflow of a typical data synthesis method, Task2Sim [76], is shown in Figure 6. Figure 7 shows samples of image data generated by the technique. The approach contrasts sharply with traditional approaches such as A3D [77] that do not make use of feedback signals from the performance on the downstream task to optimize simulation parameters. The synthesis process in the latter case, thus, involves extensive manual tuning.
@@ 3.3.2 Tabular data synthesis
Some of the most popular traditional methods for generating synthetic tabular data involve random oversampling. This approach is particularly useful for balancing imbalanced data, where minority classes are typically sampled and interpolated to create new datapoints. SMOTE [95], a seminal work among this class of approaches, is commonly used to generate synthetic data to populate tables by means of k-Nearest Neighbor algorithm which interpolates between existing data points to create new samples. These oversampling methods can be incorporated within larger neural frameworks to automate the data synthesis process . For instance, Wang and Pai [96] propose an approach that employs SMOTE together with a GAN model to generate novel tabular clinical data. Similarly, DeepSMOTE [97] combines SMOTE with an encoder-decoder sub-model for synthetic data generation. Aragao et al. [98] propose a self-balancing, synthetic tabular data generation pipeline that utilizes both undersampling and oversampling techniques within an AutoML framework to generate new data. Some new approaches bypass the step of modeling data analytically and instead learn the generation process end-to-end from data. Rashidi et al. [99], for example, devise a so-called Synthetic Tabular Neural Generator (STNG) that trains a dedicated neural network within a larger AutoML framework to specifically generate rich tabular data.
@@ 4.1 Fixed augmentation operations
Many AutoML techniques (e.g., [17, 34, 107, 30, 108, 36, 31, 109]) employ a fixed set of pre-defined transformation operations to perform data augmnetation. In these works, the transformation operations themselves are fixed and can only be varied by learning tunable parameters that control hyperparameters such as augmentation intensities and application probabilities. Thus, the task of the AutoML process in this case is reduced to simply learning the most useful combination of these fixed augmentations along with the optimal settings for these hyperparameters. In the procedure, the choice of specific transformation operations and their corresponding magnitudes to include in the search space is based on domain knowledge and intuition. Fig. 9 shows the common methods typically employed by approaches that utilize pre-defined transformation operations to perform data augmentation.
@@
The process of composing predefined augmentation operations is typically sub-divided into three main sub-problems: (1) composition of transformation operations to be used for augmentation, (2) a specification of (mostly) discrete probabilities with which the transformation operations would be applied on the input data, and (3) a specification of the applicable transformation magnitudes (i.e., augmentation intensities). Again, for simplicity, transformation magnitudes are often discrete values. The original work, AutoAugment (AA) [17], for example, uses a set of 16 common image transformation operations to augment image data for classification tasks. It trains a deep CNN model on each set of transformation operations to learn optimal parameters for effective image augmentations using validation data. Many subsequent works (e.g., [31, 30, 108, 35, 110]) adopt a similar formulation of augmentation operations in the search space as in AA [17] but utilize more efficient optimization methods and additional mechanisms to reduce computational requirements. For instance, AdvAA [110] maintained the search space specification in AA [17] but, instead of optimizing the augmentation policy hyperparameters and network parameters separately, jointly optimizes both tasks through adversarial training. Also, instead of RL, a gradient approximation method is used for the policy search. The visual results of applying fixed transformation operations with different magnitudes on image data are shown in Fig. 10. The sample results indicated are specifically for augmentation policies of FasterAA [30].
Figure 10: Augmentation policies of FasterAA [30] and the result of applying these policies on sample images of the CIFAR-10 [111] and SVHN [112] datasets.